OpenStudy (austinl):

A long straight wire has fixed negative charge with a linear charge density of magnitude 5.2 nC/m. The wire is to be enclosed by a coaxial, thin-walled, nonconducting cylindrical shell of radius 1.8 cm. The shell is to have positive charge on its outside surface with a surface charge density σ that makes the net external electric field is zero. Calculate σ.

OpenStudy (austinl):

@random231 @divu.mkr You guys have any brilliant thoughts :D

random231 (random231):

cool question working on it!

OpenStudy (anonymous):

2k (lambda) = (sigma)R/ E0 solve this and revert back

OpenStudy (anonymous):

if your answer got correct

random231 (random231):

wats k?

OpenStudy (anonymous):

@austinL you know whats k ?

OpenStudy (austinl):

k? \(8.99\times 10^9~~ Nm^2/C^2\)

OpenStudy (anonymous):

yep and k= 1/4pi E0

random231 (random231):

ok ok i know that !!!

OpenStudy (austinl):

Okay, so evidently the answer is \(45.978\times 10^{-9}~~ C/m^2\)

OpenStudy (anonymous):

and it matches with the answer of the equation?

random231 (random231):

its basically gauss's law again! see the attachment!

random231 (random231):

got it? its a bit shortcut!!!

random231 (random231):

@austinL

OpenStudy (austinl):

Yep, sorry I had some chores to do. Yep, makes sense! Thank you :D